# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
Caesar faced an important choice. On the one hand, he could obey Pompey and give up his army. But he knew the Senate didn’t like him, and he worried that he would be arrested. On the other hand, he could refuse to give up his army. But then the Senate would probably think that Caesar wanted to start a war. After thinking about it for a long time, Caesar chose the second option. He decided to attack Pompey and the Senate before they could raise an army to stop him.
Based on the passage above, how did Caesar respond to Pompey’s demands?
Choices: [(a) He disobeyed Pompey but offered to cooperate with the Senate., (b) He disobeyed Pompey and started a war against Pompey and the Senate., (c) He decided to give up his army and return to Rome to face punishment.]
Answer index: 1
Subject: Social Studies
Skill: roman-government-monarchy-to-republic
Description: According to present-day historians, the Roman Republic lasted from 509 BCE to about 27 BCE. The word republic comes from the ancient Roman term res publica. Look at the meaning of the term res publica. Then answer the question below.
res publica: a community committed to following the same rules and protecting each other’s things and families for the good of all
Based on the meaning above, which of the following statements describe something similar to what Romans meant by res publica?
Choices: [(a) a group of friends that do whatever they want, whenever they want, (b) a class project group where everyone follows the same guidelines and benefits equally from working together, (c) a soccer team where bad players have to do what the coach says, but the best players don’t]
Answer index: 1
Subject: Social Studies
Skill: french-and-english-expeditions-part-2
Description: After John Cabot’s expedition, the English government did not sponsor another expedition to the Americas for more than 50 years. Then, in the mid-1500s, the English government began sending expeditions again. Read the passage about the events in England that created new interest in expeditions. Then answer the question below.
In the middle of the 1500s, England’s main source of wealth was trading English cloth in foreign markets. Thousands of people in England made their livings by producing or selling cloth, and the English government made money by taxing the sale of cloth. But in 1551, some foreign powers stopped allowing people to purchase English cloth. Many English people who worked in the cloth industry lost their jobs, and the government could not raise as much money.
These conditions created a national crisis. The Queen of England, Elizabeth I, needed to find new sources of wealth to address the crisis. Queen Elizabeth thought that sending expeditions to the Americas might be one way to help.
Based on the passage, why did Queen Elizabeth I decide to begin sending expeditions to the Americas?
Choices: [(a) She wanted to stop Spain from purchasing English cloth., (b) She needed to find a new way to make money for England after the cloth crisis., (c) She was curious to learn more about what the Americas looked like.]
Answer index: 1
Subject: Social Studies
Skill: comparing-athens-and-sparta-part
@@
Description: Although city-states such as Athens could trade for food, Sparta was surrounded by mountains that made trade difficult. So, the Spartans relied on agriculture. The ancient Greek historian Pausanias described the role of helots, or enslaved people, in Spartan agriculture. Read the passage. Then complete the text below.
The helots had to bring half of all the crops they grew to the Spartans. Helots who disobeyed were severely punished. The poet Tyrtaeus compared the helots to donkeys exhausted by a heavy load, under painful duty to bring their Spartan masters half of the crops from the land they ploughed and cared for.
Adapted from Pausanias, Description of Greece, translated by W.H.S. Jones and H.A. Ormerod.
Complete the text.
The helots were enslaved people in Sparta. The Spartans relied on the helots ___.
Choices: [(a) to educate their children, (b) to farm their land and raise crops, (c) to trade what was farmed, (d) to train their children to farm]
Answer index: 1
Subject: Social Studies
Skill: thomas-edison
Description: Several inventors worked on lightbulbs before Thomas Edison did. How were Thomas Edison’s lightbulbs different?
@@
Choices: [(a) His lightbulbs were free., (b) His lightbulbs were made from all recycled materials., (c) His lightbulbs could last for several weeks or months., (d) His lightbulbs came in different colors.]
Answer index: 2
Subject: Social Studies
Skill: the-revolutionary-war-turning-the-tide-of-war
Description: Both the Continental Army and people on the homefront had trouble paying for the things that they needed during the war. Why?
@@
Choices: [(a) Inflation caused prices to rise., (b) The Continental Congress stopped printing Continentals., (c) The British took gold reserves from the Thirteen Colonies., (d) Most of the money was spent on military supplies.]
Answer index: 0
@@
Subject: Social Studies
Skill: foundations-of-maya-civilization
Description: Complete the text.
@@
Archaeologists believe that the ancestors of the Maya first arrived in Central America more than 12,000 years ago. At first, these people were ___ and traveled long distances to find food. But over time, they learned to grow crops instead. By at least 1500 BCE, they began to settle in permanent ___. This was the beginning of the Maya civilization.
Choices: [(a) explorers . . . cities, (b) farmers . . . palaces, (c) hunter-gatherers . . . villages]
Answer index: 2
Subject: Social Studies
Skill: classical-athens-geography-and-society
Description: Complete the text.
For most of its history, ancient Athens ruled itself and was not part of a larger country. In other words, Athens was a ___.
Choices: [(a) region, (b) city-state, (c) kingdom]
Answer index: 1
Subject: Social Studies
Skill: the-mongol-empire
Description: The founder of the Mongol Empire was a man named Temujin. He was born in Mongolia in the 1100s. The passage below comes from a Mongolian history text. In the passage, an older person is telling Temujin what Mongolia was like before he was born. Read the passage. Then answer the question below.
Before you were born, the stars turned in the heavens. Everyone was fighting. Instead of sleeping, they robbed each other of their possessions. . . . Instead of resting, they fought each other. In such a world one did not live as one wished, but rather in constant conflict. There was no rest, only battle. There was no affection, only mutual slaughter.
Adapted from The Secret History of the Mongols, paragraph 254
Based on the passage, which choice best summarizes what Mongolian society was like in the 1100s?
Choices: [(a) Mongolian society was peaceful, but foreign invaders often attacked Mongolia. As a result, Mongols were often forced to fight, and many Mongols died., (b) People in Mongolia often fought wars with China, their neighbor. As a result, people often went to China to fight., (c) People in Mongolia were fighting with one another and stealing each others’ goods. As a result, people could not live peaceful lives.]
Answer index: 2
Subject: Social Studies
Skill: foundations-of-aztec-civilization
Description: Today, the part of Central America where the Aztec Empire existed is known as Mesoamerica. There are two ways of defining a region like Mesoamerica:
using a geographic definition, which focuses on an area of land and its features, such as mountain ranges and bodies of water
using a cultural definition, which focuses on the customs and practices that people living there share
Which of the following statements uses a geographic definition of Mesoamerica?
Choices: [(a) Mesoamerica is a place where people used a 260-day calendar., (b) Mesoamerica is a place in Central America where people built religious temples in the form of pyramids., (c) Mesoamerica is a region of Central America that includes the Valley of Mexico and the Yucatan Peninsula.]
Answer index: 2
Subject: Language Arts
@@
Skill: choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
Description: Pick the uppercase letter that matches.
k
Choices: [(a) T, (b) R, (c) K]
Answer index: 2
Subject: Language Arts
@@
Skill: choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
Description: Pick the lowercase letter that matches.
C
Choices: [(a) c, (b) v, (c) o]
Answer index: 0
Subject: Language Arts
@@
Skill: choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r
Description: Pick the uppercase letter that matches. a
Choices: [(a) N, (b) A]
Answer index: 1
Subject: Language Arts
@@
Skill: choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r
@@
Description: Pick the lowercase letter that matches. A
Choices: [(a) u, (b) a, (c) n]
Answer index: 1
Subject: Language Arts
Skill: choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y
Description: Pick the lowercase letter that matches.
L
Choices: [(a) l, (b) k, (c) m]
Answer index: 0
Subject: Language Arts
@@
Skill: choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y
Description: Pick the uppercase letter that matches.
m
Choices: [(a) H, (b) J, (c) M]
Answer index: 2
Subject: Language Arts
Skill: distinguish-characters-points-of-view
Description: The following texts both describe a school bus ride on the first day of school.
This text is told from the point of view of Jeffrey, a student going to a new school:
@@
I felt ready for my first shift as a school bus driver. I turned the long bus at each corner with ease. My cheerful mood made the students smile. I was proud of myself as the bus approached the brick school building. My first trip had gone well!
How is Ms. Romero’s point of view different from Jeffrey’s?
Choices: [(a) Ms. Romero describes feeling uneasy about a new experience., (b) Ms. Romero expresses a feeling of pride about the first day of school.]
Answer index: 1
Subject: Language Arts
Skill: determine-the-meanings-of-words-with-greek-and-latin-roots
Description: What does the word autopilot mean?
@@
Choices: [(a) a tool that pilots use when they want to make the airplane fly faster, (b) a device that flies an airplane on its own, without a pilot, (c) a map or drawing that shows an airplane’s route]
Answer index: 1
Subject: Language Arts
Skill: describe-the-difference-between-related-words
Description: What is the difference between being vicious and being mean?
Choices: [(a) being vicious is worse, (b) being mean shows more anger]
Answer index: 0
Subject: Language Arts
Skill: which-word-does-not-rhyme
Description: Which word does not rhyme?
Choices: [(a) pack, (b) pick, (c) sick]
Answer index: 0
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-sight-word
Description: Complete the sentence.
___ pig is big.
Choices: [(a) That, (b) Went]
Answer index: 0
Subject: Language Arts
Skill: to-have-use-the-correct-form
Description: Complete the sentence.
Mrs. Barton’s rug ___ blue flowers on it.
Choices: [(a) has, (b) have]
Answer index: 0
Subject: Language Arts
Skill: complete-the-sentence-with-a-two-syllable-word
Description: Which word makes more sense in the sentence?
We ride in the ___
Choices: [(a) wagon, (b) robin]
Answer index: 0
Subject: Language Arts
Skill: complete-the-sentence-with-a-three-letter-consonant-blend-word
Description: Which word makes more sense in the sentence?
Look, that big fish is swimming up the ___
Choices: [(a) stream, (b) strict]
Answer index: 0
Subject: Language Arts
Skill: use-subordinating-conjunctions
Description: Complete the sentence with the better conjunction.
___ Joseph arrives soon, we won’t get to the movie on time.
Choices: [(a) Unless, (b) After]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-homophone
Description: Complete the sentence with the correct homophone.
@@
Given that there were more than one hundred guests at the wedding, how many ___ do you think Martina and Cooper have received?
Choices: [(a) presence, (b) presents]
Answer index: 1
Subject: Language Arts
Skill: use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
Description: The word millimeter contains the root mille. What does the root mille mean?
Choices: [(a) measure, (b) one hundred, (c) one thousand]
Answer index: 2
Subject: Language Arts
Skill: analogies
Description: Complete the analogy.
Ohio is to United States as Earth is to
Choices: [(a) moon, (b) solar system]
Answer index: 1
@@
Subject: Language Arts
Skill: is-it-a-complete-sentence-or-a-fragment
Description: Is this a complete sentence?
Monica speaks four languages.
Choices: [(a) Monica speaks four languages.]
Answer index: 0
Subject: Language Arts
Skill: is-it-a-phrase-or-a-clause
Description: Is the group of words in bold a phrase or a clause?
*Emily’s suitcase was too large to fit in the overhead bin**, so unfortunately she had to check it.
Choices: [(a) phrase, (b) clause]
Answer index: 1
Subject: Language Arts
Skill: choose-between-subject-and-object-pronouns
Description: Select the correct pronoun to complete the sentence.
@@
"Thanks to your efforts, ___ have emerged as leaders in the energy industry," Mr. Bullock announced at the company meeting.
Choices: [(a) us, (b) we]
Answer index: 1
Subject: Language Arts
Skill: is-the-sentence-simple-compound-complex-or-compound-complex
Description: Which type of sentence is this?
@@
Henry always approaches difficult tasks enthusiastically, and he frequently motivates others with his energy and fervor.
Choices: [(a) simple, (b) compound, (c) complex, (d) compound-complex]
Answer index: 1
Subject: Language Arts
Skill: analogies-challenge
Description: Complete the analogy.
teacher is to instructor as doctor is to
Choices: [(a) profession, (b) physician, (c) medicine]
Answer index: 1
Subject: Language Arts
Skill: use-time-order-words
Description: Complete the text with the better word.
, Jerry jogs for a mile. Then, he rests for ten minutes.
Choices: [(a) First, (b) Finally]
Answer index: 0
Subject: Language Arts
Skill: choose-reasons-to-support-an-opinion
Description: Read the opinion below.
Swamps should be protected and valued.
Select the reason that best supports this opinion.
@@
Choices: [(a) Swamps do an excellent job of cleaning an area’s water supply., (b) Many dangerous creatures, such as alligators and poisonous snakes, live in swamps.]
Answer index: 0
Subject: Language Arts
Skill: is-it-a-complete-sentence-or-a-run-on
Description: Which is a **run-on sentence**?
@@
Choices: [(a) The heart is a symbol of love., (b) She went to the tailor, he sewed her pants.]
Answer index: 1
Subject: Language Arts
Skill: read-about-famous-people
Description: Review the text.
title: Becoming Amazing
Copeland’s success made her a role model for others, too. She was on the cover of Time magazine as one of the "100 Most Influential People" in 2015. She wrote a best-selling book about her life, called Life in Motion: An Unlikely Ballerina. She also wrote a children’s book, Firebird, in which she tells others like herself that they, too, can become amazing.
Which is the best summary of the text?
@@
Choices: [(a) Misty Copeland beat many challenges to become a famous ballet dancer. She started later than most dancers and suffered broken bones. But with the help of a coach and a ballet teacher, she became the first African American principal dancer for ABT. She also became an author and role model., (b) Misty Copeland never stopped working to become a famous ballet dancer. She started training very young and became an amazing dancer by age thirteen. After becoming the first African American woman dancer for ABT, she injured herself badly, so she became a model and an author instead., (c) Misty Copeland started dancing at age thirteen, much later than most ballet dancers. She was encouraged by her middle-school drill team coach and a ballet teacher at the Boys & Girls Club. With a lot of hard work, she became a professional dancer with ABT at the age of eighteen.]
Answer index: 0
Subject: Language Arts
Skill: choose-the-synonym
Description: Which word is a synonym of **bag**?
Choices: [(a) sack, (b) contain]
Answer index: 0
Subject: Language Arts
Skill: identify-story-elements
Description: Review the story.
title: Contest on Mars Colony
He could see the kids taking deep breaths of the forest air. Whether or not he won, their smiles were prize enough for him.
What is the **solution** to the problem in the story?
@@
Choices: [(a) Axel comes up with a high-tech app idea., (b) Zane invents a new alarm that will play beautiful music instead of buzzing., (c) Axel realizes he can improve the Colony by growing a forest., (d) Axel’s father decides they’ll move back to Earth.]
Answer index: 2
Subject: Language Arts
Skill: use-academic-vocabulary-in-context
Description: Look at the word **residence** in the passage below.
@@
In many homes, basements are used for ordinary activities, such as washing clothes. The U.S. president’s basement is much more interesting. In the White House, the president’s **residence**, the basement contains a woodshop and a dentist’s office. And when the president’s family wants to have some fun, they can go bowling in their own basement bowling alley!
What is the meaning of **residence** as used in the passage?
Choices: [(a) a place where someone lives, (b) a place where parties are held, (c) a place that people like to visit]
Answer index: 0
Subject: Language Arts
Skill: use-adjectives-to-compare
Description: Complete the sentence with the correct form of the adjective.
The traffic on Main Street is much ___ than the traffic on the highway.
Choices: [(a) slower, (b) slowest]
Answer index: 0
Subject: Language Arts
Skill: select-the-words-that-dont-belong
Description: Select the place that "doesn’t" belong.
Choices: [(a) library, (b) meadow, (c) forest, (d) field]
Answer index: 0
Subject: Language Arts
Skill: capitalizing-titles
Description: Which correctly shows the title of a poem?
Choices: [(a) "The Mountain and the Squirrel", (b) "the Mountain and the Squirrel"]
Answer index: 0
Subject: Language Arts
Skill: read-historical-fiction
Description: Read the story.
title: Strangers on My Shore
@@
Many changes have come in the 250 years that have passed since that day. Now, big cities dot my landscape, and two-legged creatures fly from around the world to visit me. Though many now call me New Zealand, I am still the same island, shimmering green under the clouds. After all these years, my waters and my mountains remain, and my bays and rivers endure.
Based on paragraphs 1 through 3, how does the fact that the narrator is an island rather than a person affect this telling of New Zealand’s history?
Choices: [(a) It allows the narrator to compare the landforms in New Zealand to other places in the Pacific., (b) It allows the story to start long before the arrival of humans, with the formation of the land., (c) It allows the story to focus only on the geography of New Zealand, not the people who live there., (d) It allows the narrator to refer to the island by the present-day name of "New Zealand."]
Answer index: 1

### v2

@@ Appendix D Summary of Skills
Subject: Social Studies Skill: caesar-and-the-end-of-the-roman-republic Description: Pompey agreed to help the Senate and try to prevent Caesar from becoming more powerful. Read the passage. Then answer the question below. In January of 49 BCE, Caesar was with his army far away from Rome. Pompey wrote to Caesar and demanded that Caesar give up his army and return to Rome. Caesar faced an important choice. On the one hand, he could obey Pompey and give up his army. But he knew the Senate didn’t like him, and he worried that he would be arrested. On the other hand, he could refuse to give up his army. But then the Senate would probably think that Caesar wanted to start a war. After thinking about it for a long time, Caesar chose the second option. He decided to attack Pompey and the Senate before they could raise an army to stop him. Based on the passage above, how did Caesar respond to Pompey’s demands? Choices: [(a) He disobeyed Pompey but offered to cooperate with the Senate., (b) He disobeyed Pompey and started a war against Pompey and the Senate., (c) He decided to give up his army and return to Rome to face punishment.] Answer index: 1
Subject: Social Studies Skill: roman-government-monarchy-to-republic Description: According to present-day historians, the Roman Republic lasted from 509 BCE to about 27 BCE. The word republic comes from the ancient Roman term res publica. Look at the meaning of the term res publica. Then answer the question below. res publica: a community committed to following the same rules and protecting each other’s things and families for the good of all Based on the meaning above, which of the following statements describe something similar to what Romans meant by res publica? Choices: [(a) a group of friends that do whatever they want, whenever they want, (b) a class project group where everyone follows the same guidelines and benefits equally from working together, (c) a soccer team where bad players have to do what the coach says, but the best players don’t] Answer index: 1
Subject: Social Studies Skill: french-and-english-expeditions-part-2 Description: After John Cabot’s expedition, the English government did not sponsor another expedition to the Americas for more than 50 years. Then, in the mid-1500s, the English government began sending expeditions again. Read the passage about the events in England that created new interest in expeditions. Then answer the question below. In the middle of the 1500s, England’s main source of wealth was trading English cloth in foreign markets. Thousands of people in England made their livings by producing or selling cloth, and the English government made money by taxing the sale of cloth. But in 1551, some foreign powers stopped allowing people to purchase English cloth. Many English people who worked in the cloth industry lost their jobs, and the government could not raise as much money. These conditions created a national crisis. The Queen of England, Elizabeth I, needed to find new sources of wealth to address the crisis. Queen Elizabeth thought that sending expeditions to the Americas might be one way to help. Based on the passage, why did Queen Elizabeth I decide to begin sending expeditions to the Americas? Choices: [(a) She wanted to stop Spain from purchasing English cloth., (b) She needed to find a new way to make money for England after the cloth crisis., (c) She was curious to learn more about what the Americas looked like.] Answer index: 1
@@
Subject: Social Studies Skill: comparing-athens-and-sparta-part Description: Although city-states such as Athens could trade for food, Sparta was surrounded by mountains that made trade difficult. So, the Spartans relied on agriculture. The ancient Greek historian Pausanias described the role of helots, or enslaved people, in Spartan agriculture. Read the passage. Then complete the text below. The helots had to bring half of all the crops they grew to the Spartans. Helots who disobeyed were severely punished. The poet Tyrtaeus compared the helots to donkeys exhausted by a heavy load, under painful duty to bring their Spartan masters half of the crops from the land they ploughed and cared for. Adapted from Pausanias, Description of Greece, translated by W.H.S. Jones and H.A. Ormerod. Complete the text. The helots were enslaved people in Sparta. The Spartans relied on the helots ___. Choices: [(a) to educate their children, (b) to farm their land and raise crops, (c) to trade what was farmed, (d) to train their children to farm] Answer index: 1
@@
Subject: Social Studies Skill: thomas-edison Description: Several inventors worked on lightbulbs before Thomas Edison did. How were Thomas Edison’s lightbulbs different? Choices: [(a) His lightbulbs were free., (b) His lightbulbs were made from all recycled materials., (c) His lightbulbs could last for several weeks or months., (d) His lightbulbs came in different colors.] Answer index: 2
@@
Subject: Social Studies Skill: the-revolutionary-war-turning-the-tide-of-war Description: Both the Continental Army and people on the homefront had trouble paying for the things that they needed during the war. Why? Choices: [(a) Inflation caused prices to rise., (b) The Continental Congress stopped printing Continentals., (c) The British took gold reserves from the Thirteen Colonies., (d) Most of the money was spent on military supplies.] Answer index: 0
@@
Subject: Social Studies Skill: foundations-of-maya-civilization Description: Complete the text. Archaeologists believe that the ancestors of the Maya first arrived in Central America more than 12,000 years ago. At first, these people were ___ and traveled long distances to find food. But over time, they learned to grow crops instead. By at least 1500 BCE, they began to settle in permanent ___. This was the beginning of the Maya civilization. Choices: [(a) explorers . . . cities, (b) farmers . . . palaces, (c) hunter-gatherers . . . villages] Answer index: 2
Subject: Social Studies Skill: classical-athens-geography-and-society Description: Complete the text. For most of its history, ancient Athens ruled itself and was not part of a larger country. In other words, Athens was a ___. Choices: [(a) region, (b) city-state, (c) kingdom] Answer index: 1
Subject: Social Studies Skill: the-mongol-empire Description: The founder of the Mongol Empire was a man named Temujin. He was born in Mongolia in the 1100s. The passage below comes from a Mongolian history text. In the passage, an older person is telling Temujin what Mongolia was like before he was born. Read the passage. Then answer the question below. Before you were born, the stars turned in the heavens. Everyone was fighting. Instead of sleeping, they robbed each other of their possessions. . . . Instead of resting, they fought each other. In such a world one did not live as one wished, but rather in constant conflict. There was no rest, only battle. There was no affection, only mutual slaughter. Adapted from The Secret History of the Mongols, paragraph 254 Based on the passage, which choice best summarizes what Mongolian society was like in the 1100s? Choices: [(a) Mongolian society was peaceful, but foreign invaders often attacked Mongolia. As a result, Mongols were often forced to fight, and many Mongols died., (b) People in Mongolia often fought wars with China, their neighbor. As a result, people often went to China to fight., (c) People in Mongolia were fighting with one another and stealing each others’ goods. As a result, people could not live peaceful lives.] Answer index: 2
Subject: Social Studies Skill: foundations-of-aztec-civilization Description: Today, the part of Central America where the Aztec Empire existed is known as Mesoamerica. There are two ways of defining a region like Mesoamerica: using a geographic definition, which focuses on an area of land and its features, such as mountain ranges and bodies of water using a cultural definition, which focuses on the customs and practices that people living there share Which of the following statements uses a geographic definition of Mesoamerica? Choices: [(a) Mesoamerica is a place where people used a 260-day calendar., (b) Mesoamerica is a place in Central America where people built religious temples in the form of pyramids., (c) Mesoamerica is a region of Central America that includes the Valley of Mexico and the Yucatan Peninsula.] Answer index: 2
@@
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z Description: Pick the uppercase letter that matches. k Choices: [(a) T, (b) R, (c) K] Answer index: 2
@@
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z Description: Pick the lowercase letter that matches. C Choices: [(a) c, (b) v, (c) o] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r Description: Pick the uppercase letter that matches. a Choices: [(a) N, (b) A] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r Description: Pick the lowercase letter that matches. A Choices: [(a) u, (b) a, (c) n] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y Description: Pick the lowercase letter that matches. L Choices: [(a) l, (b) k, (c) m] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y Description: Pick the uppercase letter that matches. m Choices: [(a) H, (b) J, (c) M] Answer index: 2
@@
Subject: Language Arts Skill: distinguish-characters-points-of-view Description: The following texts both describe a school bus ride on the first day of school. This text is told from the point of view of Jeffrey, a student going to a new school: I felt ready for my first shift as a school bus driver. I turned the long bus at each corner with ease. My cheerful mood made the students smile. I was proud of myself as the bus approached the brick school building. My first trip had gone well! How is Ms. Romero’s point of view different from Jeffrey’s? Choices: [(a) Ms. Romero describes feeling uneasy about a new experience., (b) Ms. Romero expresses a feeling of pride about the first day of school.] Answer index: 1
@@
Subject: Language Arts Skill: determine-the-meanings-of-words-with-greek-and-latin-roots Description: What does the word autopilot mean? Choices: [(a) a tool that pilots use when they want to make the airplane fly faster, (b) a device that flies an airplane on its own, without a pilot, (c) a map or drawing that shows an airplane’s route] Answer index: 1
Subject: Language Arts Skill: describe-the-difference-between-related-words Description: What is the difference between being vicious and being mean? Choices: [(a) being vicious is worse, (b) being mean shows more anger] Answer index: 0
Subject: Language Arts Skill: which-word-does-not-rhyme Description: Which word does not rhyme? Choices: [(a) pack, (b) pick, (c) sick] Answer index: 0
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-sight-word Description: Complete the sentence. ___ pig is big. Choices: [(a) That, (b) Went] Answer index: 0
Subject: Language Arts Skill: to-have-use-the-correct-form Description: Complete the sentence. Mrs. Barton’s rug ___ blue flowers on it. Choices: [(a) has, (b) have] Answer index: 0
Subject: Language Arts Skill: complete-the-sentence-with-a-two-syllable-word Description: Which word makes more sense in the sentence? We ride in the ___ Choices: [(a) wagon, (b) robin] Answer index: 0
Subject: Language Arts Skill: complete-the-sentence-with-a-three-letter-consonant-blend-word Description: Which word makes more sense in the sentence? Look, that big fish is swimming up the ___ Choices: [(a) stream, (b) strict] Answer index: 0
Subject: Language Arts Skill: use-subordinating-conjunctions Description: Complete the sentence with the better conjunction. ___ Joseph arrives soon, we won’t get to the movie on time. Choices: [(a) Unless, (b) After] Answer index: 0
@@
Subject: Language Arts Skill: use-the-correct-homophone Description: Complete the sentence with the correct homophone. Given that there were more than one hundred guests at the wedding, how many ___ do you think Martina and Cooper have received? Choices: [(a) presence, (b) presents] Answer index: 1
Subject: Language Arts Skill: use-words-as-clues-to-the-meanings-of-greek-and-latin-roots Description: The word millimeter contains the root mille. What does the root mille mean? Choices: [(a) measure, (b) one hundred, (c) one thousand] Answer index: 2
Subject: Language Arts Skill: analogies Description: Complete the analogy. Ohio is to United States as Earth is to Choices: [(a) moon, (b) solar system] Answer index: 1
@@
Subject: Language Arts Skill: is-it-a-complete-sentence-or-a-fragment Description: Is this a complete sentence? Monica speaks four languages. Choices: [(a) Monica speaks four languages.] Answer index: 0
Subject: Language Arts Skill: is-it-a-phrase-or-a-clause Description: Is the group of words in bold a phrase or a clause? **Emily’s suitcase was too large to fit in the overhead bin**, so unfortunately she had to check it. Choices: [(a) phrase, (b) clause] Answer index: 1
@@
Subject: Language Arts Skill: choose-between-subject-and-object-pronouns Description: Select the correct pronoun to complete the sentence. "Thanks to your efforts, ___ have emerged as leaders in the energy industry," Mr. Bullock announced at the company meeting. Choices: [(a) us, (b) we] Answer index: 1
@@
Subject: Language Arts Skill: is-the-sentence-simple-compound-complex-or-compound-complex Description: Which type of sentence is this? Henry always approaches difficult tasks enthusiastically, and he frequently motivates others with his energy and fervor. Choices: [(a) simple, (b) compound, (c) complex, (d) compound-complex] Answer index: 1
Subject: Language Arts Skill: analogies-challenge Description: Complete the analogy. teacher is to instructor as doctor is to Choices: [(a) profession, (b) physician, (c) medicine] Answer index: 1
Subject: Language Arts Skill: use-time-order-words Description: Complete the text with the better word. , Jerry jogs for a mile. Then, he rests for ten minutes. Choices: [(a) First, (b) Finally] Answer index: 0
@@
Subject: Language Arts Skill: choose-reasons-to-support-an-opinion Description: Read the opinion below. Swamps should be protected and valued. Select the reason that best supports this opinion. Choices: [(a) Swamps do an excellent job of cleaning an area’s water supply., (b) Many dangerous creatures, such as alligators and poisonous snakes, live in swamps.] Answer index: 0
@@
Subject: Language Arts Skill: is-it-a-complete-sentence-or-a-run-on Description: Which is a **run-on sentence**? Choices: [(a) The heart is a symbol of love., (b) She went to the tailor, he sewed her pants.] Answer index: 1
@@
Subject: Language Arts Skill: read-about-famous-people Description: Review the text. title: Becoming Amazing Copeland’s success made her a role model for others, too. She was on the cover of Time magazine as one of the "100 Most Influential People" in 2015. She wrote a best-selling book about her life, called Life in Motion: An Unlikely Ballerina. She also wrote a children’s book, Firebird, in which she tells others like herself that they, too, can become amazing. Which is the best summary of the text? Choices: [(a) Misty Copeland beat many challenges to become a famous ballet dancer. She started later than most dancers and suffered broken bones. But with the help of a coach and a ballet teacher, she became the first African American principal dancer for ABT. She also became an author and role model., (b) Misty Copeland never stopped working to become a famous ballet dancer. She started training very young and became an amazing dancer by age thirteen. After becoming the first African American woman dancer for ABT, she injured herself badly, so she became a model and an author instead., (c) Misty Copeland started dancing at age thirteen, much later than most ballet dancers. She was encouraged by her middle-school drill team coach and a ballet teacher at the Boys & Girls Club. With a lot of hard work, she became a professional dancer with ABT at the age of eighteen.] Answer index: 0
Subject: Language Arts Skill: choose-the-synonym Description: Which word is a synonym of **bag**? Choices: [(a) sack, (b) contain] Answer index: 0
@@
Subject: Language Arts Skill: identify-story-elements Description: Review the story. title: Contest on Mars Colony He could see the kids taking deep breaths of the forest air. Whether or not he won, their smiles were prize enough for him. What is the **solution** to the problem in the story? Choices: [(a) Axel comes up with a high-tech app idea., (b) Zane invents a new alarm that will play beautiful music instead of buzzing., (c) Axel realizes he can improve the Colony by growing a forest., (d) Axel’s father decides they’ll move back to Earth.] Answer index: 2
@@
Subject: Language Arts Skill: use-academic-vocabulary-in-context Description: Look at the word **residence** in the passage below. In many homes, basements are used for ordinary activities, such as washing clothes. The U.S. president’s basement is much more interesting. In the White House, the president’s **residence**, the basement contains a woodshop and a dentist’s office. And when the president’s family wants to have some fun, they can go bowling in their own basement bowling alley! What is the meaning of **residence** as used in the passage? Choices: [(a) a place where someone lives, (b) a place where parties are held, (c) a place that people like to visit] Answer index: 0
Subject: Language Arts Skill: use-adjectives-to-compare Description: Complete the sentence with the correct form of the adjective. The traffic on Main Street is much ___ than the traffic on the highway. Choices: [(a) slower, (b) slowest] Answer index: 0
Subject: Language Arts Skill: select-the-words-that-dont-belong Description: Select the place that "doesn’t" belong. Choices: [(a) library, (b) meadow, (c) forest, (d) field] Answer index: 0
Subject: Language Arts Skill: capitalizing-titles Description: Which correctly shows the title of a poem? Choices: [(a) "The Mountain and the Squirrel", (b) "the Mountain and the Squirrel"] Answer index: 0
@@
Subject: Language Arts Skill: read-historical-fiction Description: Read the story. title: Strangers on My Shore Many changes have come in the 250 years that have passed since that day. Now, big cities dot my landscape, and two-legged creatures fly from around the world to visit me. Though many now call me New Zealand, I am still the same island, shimmering green under the clouds. After all these years, my waters and my mountains remain, and my bays and rivers endure. Based on paragraphs 1 through 3, how does the fact that the narrator is an island rather than a person affect this telling of New Zealand’s history? Choices: [(a) It allows the narrator to compare the landforms in New Zealand to other places in the Pacific., (b) It allows the story to start long before the arrival of humans, with the formation of the land., (c) It allows the story to focus only on the geography of New Zealand, not the people who live there., (d) It allows the narrator to refer to the island by the present-day name of "New Zealand."] Answer index: 1